I'm stuck on this problem. I need it to be solved for me step by step.
Four oranges and five apples cost $3.56. Four oranges and three apples cost $2.76. Find the cost of an orange and the cost of an apple.

Answer by Abbey(339)

Let the cost of oranges be x
Let the cost of apples be y
Then:
4x+5Y=3.56
4x+3Y=2.76
subtract the second equation from the first
2y=.80
divide both sides by 2
y=.40
put this back into either of the original equations to find out how much an orange costs
4x+5(.40)=3.56
4x+2.00=3.56
4x=1.56
x=.39 cents
check this answer by using the other equation:
4(.39)+3(.40)=2.76
1.56+1.20=2.76
This is true, so oranges cost .39 cents each and apples cost .40 cents each.